OpenStudy (anonymous):

Could someone not just tell me the answer, but show me how to do this problem? Use the double angle formulas to help you find all solutions to the following on the interval [0, 2pi). sin(2x) + cosx = 0

terenzreignz (terenzreignz):

Gladly :3 Do you know the double angle identity for sines?\[\large \sin(2x) = \color{red}?\]

OpenStudy (anonymous):

2sinxcosx?

terenzreignz (terenzreignz):

That's right :) \[\Large 2\sin(x)\cos(x) +\cos(x) = 0\] Get it so far? :)

OpenStudy (anonymous):

Yeah.

terenzreignz (terenzreignz):

Well, this works much like solving algebraic equations. You can factor our the cos(x), giving you... \[\Large \color{red}{\cos(x)}\color{blue}{[2\sin(x)+1]}=0\]

OpenStudy (anonymous):

Ok.

terenzreignz (terenzreignz):

Well, remember, when you have a pair of factors that equal zero.. ab = 0 This can only happen when at least one of the factors is zero. So...either a = 0 or b = 0

OpenStudy (anonymous):

Ah! So cosx=0 and 2sinx=-1 right?

terenzreignz (terenzreignz):

Very good. But use an 'or' and not an 'and' :) Because I doubt those two may be true at the same time :D

terenzreignz (terenzreignz):

cos(x) = 0 What are its solutions from \(\large [0, 2\pi)\)

OpenStudy (anonymous):

How do you check if the values are true or not?

terenzreignz (terenzreignz):

We will try it later :) What values of x would make the equation cos(x) = 0 true? (Assuming x is in \([0, 2\pi)\) )

OpenStudy (anonymous):

pi/2 right?

terenzreignz (terenzreignz):

And there's another one :)

OpenStudy (anonymous):

3pi/2

terenzreignz (terenzreignz):

okay, good :) Now what about the values that would make 2sin(x) = -1 true? This is essentially the same as \[\Large \sin(x) = -\frac 12\]

OpenStudy (anonymous):

7pi/6 and 11pi/6

terenzreignz (terenzreignz):

:) And you have now found.. the solutions. You want to test them out? :)

OpenStudy (anonymous):

Yes please.

terenzreignz (terenzreignz):

Our solution set is... \[\Large \left\{\frac \pi 2 , \frac{7\pi}6 , \frac{3\pi}2, \frac{11\pi}6\right\}\]

terenzreignz (terenzreignz):

Originally the equation was \[\Large \sin(2x) + \cos(x) = 0\] Let's try \(\large \frac \pi 2\) shall we? \[\Large \sin\left(2\cdot \frac \pi 2\right) + \cos\left(\frac \pi 2\right)\] Right?

OpenStudy (anonymous):

Yeah.

terenzreignz (terenzreignz):

Well... 2 cancels out here...\[\Large \sin\left(\cancel2\cdot \frac \pi {\cancel2}\right) + \cos\left(\frac \pi 2\right)\] Leaving \[\Large \sin\left(\pi\right) + \cos\left(\frac \pi 2\right)\]

OpenStudy (anonymous):

That is zero.

terenzreignz (terenzreignz):

mhmm :)

terenzreignz (terenzreignz):

Let's test \(\Large \frac{7\pi}{6}\)

OpenStudy (anonymous):

So it is true.

OpenStudy (anonymous):

Sorry, internet lag.

terenzreignz (terenzreignz):

It's true for \(\Large \frac \pi 2\) You can take my word for it and accept that it'd also be true for all the others in our little solution set :D

terenzreignz (terenzreignz):

Or we could also test them out ^_^

terenzreignz (terenzreignz):

\[\Large \sin\left(2\cdot \frac{7\pi}6\right) + \cos\left(\frac {7\pi} {6}\right)\]

terenzreignz (terenzreignz):

\[\Large \sin\left(\cancel2\cdot \frac{7\pi}{\cancel6_3}\right) + \cos\left(\frac {7\pi} {6}\right)\]

terenzreignz (terenzreignz):

\[\Large \sin\left(\frac{7\pi}3\right) + \cos\left(\frac {7\pi} {6}\right)\]

OpenStudy (anonymous):

That is zero also. Sorry, lagging

terenzreignz (terenzreignz):

Are you sure? :)

OpenStudy (anonymous):

I think so...

terenzreignz (terenzreignz):

If I'm not mistaken, this results in \[\Large \frac{\sqrt 3}2 - \frac{\sqrt 3}2\] So yeah :D

terenzreignz (terenzreignz):

Wanna try \(\Large \frac{3\pi}2\) ? :)

OpenStudy (anonymous):

Ok.

terenzreignz (terenzreignz):

\[\Large \sin\left(2\cdot \frac{3\pi}2\right) + \cos\left(\frac {3\pi} {2}\right)\]

OpenStudy (anonymous):

I did the others, they are also both zero, right?

terenzreignz (terenzreignz):

Well, yes :D

OpenStudy (anonymous):

Could you help just start one of the half angle problems that I have?

terenzreignz (terenzreignz):

How specific :D Okay, go ^_^

OpenStudy (anonymous):

cos^2(x)=sin^2(x/2)

terenzreignz (terenzreignz):

Okay, so... what is the value for \[\Large \sin\left(\frac x 2\right)=\color{red}?\]

OpenStudy (anonymous):

+-sqrt((1-cosa)/2)

terenzreignz (terenzreignz):

x. not a :P \[\Large \sin\left(\frac{x}2\right) = \pm \sqrt{\frac{1-\cos(x)}2}\]

terenzreignz (terenzreignz):

And good, by the way :) Thankfully, in this specific problem that \(\large \pm\) sign is rather irrelevant. Can you see why? ^_^

terenzreignz (terenzreignz):

\[\huge \cos^2(x) = \sin^2\left(\frac{x}{2}\right)\]

OpenStudy (anonymous):

We know it has to be positive because of the starting equation before?

terenzreignz (terenzreignz):

It is irrelevant because of this bit \[\huge \cos^2(x) = \sin^{\boxed {\color{red}{2}}}\left(\frac{x}{2}\right)\] That square would just remove the radical, regardless of whether or not it was positive :)

OpenStudy (anonymous):

Ah, I see.

terenzreignz (terenzreignz):

\[\Large \sin\left(\frac{x}2\right) = \pm \sqrt{\frac{1-\cos(x)}2}\] Can only mean that \[\Large \sin^{\color{red}2}\left(\frac{x}2\right) ={\frac{1-\cos(x)}2}\] Right? :)

OpenStudy (anonymous):

Oh, so the square root cancels out.

terenzreignz (terenzreignz):

Yup :) Leaving you with \[\Large \cos^2(x) = \frac{1-\cos(x)}2\] If you like, you can let u = cos(x) So it now becomes \[\Large u^2 = \frac{1-u}2\]

terenzreignz (terenzreignz):

And you can solve for u (after some manipulation) using the quadratic formula, or any method which makes you comfy :3

OpenStudy (anonymous):

Ok, well thanks!

terenzreignz (terenzreignz):

No problem :) After solving for u, change it back to cos(x) again and like what we did, look for values of x which would make cos(x) = <your solution> true. ^_^

OpenStudy (anonymous):

You should be proud. Not everyone takes a big portion of their time just to help people :)